# Explainable Graph Neural Networks Under Fire

### v1

@@ 3.2 Relaxations and Loss Definitions
Relaxations. We assume that edge flipping graph $\mathcal{M}$ corresponding to $\mathbf{P}\in\{0,1\}^{n\times n}$ (namely using $\mathbf{P}$ as the adjacency matrix) is a Gilbert random graph. In other words, edge occurrences are conditionally independent from each other. As a result, with slight abuse of notations, the probability of graph $\mathcal{M}$ can be factorized as $P(\mathcal{M})=\prod_{1\leq i,j\leq n}P(m_{ij})$, dubbed $\tilde{\mathbf{P}}$. Next, we instantiate $P(m_{ij})$ using a Bernoulli distribution, namely $P(m_{ij})\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{ij})$ with $P(m_{ij}=1)=\tilde{\mathbf{P}}_{ij}$. As a result, Eq 7 is equivalent to
@@
In this way, we relax the $\mathbf{P}\in\{0,1\}^{n\times n}$ from binary variables to continuous variables $\tilde{\mathbf{P}}\in[0,1]^{n\times n}$. Hence, we can utilise gradient-based methods to optimize the objective function. At the last step of the attack, we then discretize the result (see Section 3.3).
@@
$\underset{\tilde{\mathbf{P}}}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X})):=\underset{\tilde{\mathbf{P}}}{\max}~{}H(f_{\theta}(\mathbf{A},\mathbf{X}))-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))\$ (9)
$\propto\underset{\tilde{\mathbf{P}}}{\max}~{}-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))\approx\underset{\tilde{\mathbf{P}}}{\max}~{}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}})).$ (10)
Eq 9 holds by definition of mutual information; Eq 10 left half holds as $\theta,\mathbf{A},\mathbf{X}$ are fixed; Eq 10 right half holds when we replace conditional entropy $H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ with cross entropy $H(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$. We define $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}+\tilde{\mathbf{P}}_{ij}$ if $\mathbf{A}_{ij}=0$, and $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}-\tilde{\mathbf{P}}_{ij}$ otherwise.
@@
$\underset{\tilde{\mathbf{P}}}{\min}~{}\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X})))$ $\propto\underset{\tilde{\mathbf{P}}}{\max}~{}\text{Dist}(\mathbf{M}_{s},\hat{\mathbf{M}}_{s})$ (11)
@@
Eq 11 holds by definition; Eq 12 holds when we define the $Dist$ function as the cosine distance metric, where $\text{vec}(\cdot)$ flattens a matrix to a vector, and $\|\cdot\|_{F}$ represents the Frobenius Norm.
Final Relaxed Objective. As a result, optimizing Eq 7 amounts to optimizing
$\underset{\tilde{\mathbf{P}}\in[0,1]^{n\times n}\text{~{}s.t.~{}}\sum\tilde{\mathbf{P}}_{ij}<B}{\max}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}}))+\beta*\left(1-\frac{\text{vec}(\mathbf{M}_{pr})\cdot\text{vec}(\hat{\mathbf{M}}_{pr})}{\|\mathbf{M}_{pr}\|_{F}\|\hat{\mathbf{M}}_{pr}\|_{F}}\right),$ (13)
where $\mathcal{L}(\tilde{\mathbf{P}})$ is differentiable w.r.t. $\tilde{\mathbf{P}}$.
3.3 Optimization
We employ Projected Gradient Ascent Algorithm (based on Xu et al. (2019)) to optimize objective function (13), and the details are given in Algorithm 1. To go back from the continuous $\tilde{\mathbf{P}}_{ij}$ to discrete ${\mathbf{P}}_{ij}$, we perform Bernoulli sampling with ${\mathbf{P}}_{ij}\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{ij})$ after optimization (Line 10). We note that the Projected Randomized Block Coordinate Descent (PR-BCD) (Geisler et al., 2021) can perform this optimization more efficiently, but do not compare both methods due to the low scalability of the evaluated explainers. The complexity analysis of GXAttack is given in Appendix C.
@@ 3.3 Optimization
10:    $\tilde{\mathbf{P}}_{t}\leftarrow\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})\leq B]}(\tilde{\mathbf{P}}_{t})$triangleright Project gradients, see Appendix B for more detail
@@
4 Experimental Setup
@@ 4.1 Datasets
To answer these questions, we need to evaluate the GNN explanation quality before and after the attacks. However, as pointed out by Agarwal et al. (2023), evaluating the quality of post-hoc GNN explanations is challenging. This is because existing benchmark graph datasets lack ground-truth explanations or their explanations are not reliable. To mitigate this, they proposed a synthetic graph data generator ShapeGGen. It can generate a variety of graph datasets with ground-truth explanations, where graph size, node degree distributions, the ground-truth explanation sizes, and more can be varied. As a result, ShapeGGen allows us to mimic graph data in various real-world applications. More importantly, ShapeGGen can ensure that the generated data and its ground-truth explanations do not suffer from GNN explanation evaluation pitfalls (Faber et al., 2021), namely trivial explanations, redundant explanations, and weak GNN predictors. This makes ShapeGGen very suitable for studying the limitations of GNN explainers. We consider the synthetic datasets in Table 3 (See Appendix D).
@@
Why not real-world datasets? We do not consider real-world datasets for two key reasons: 1) synthetic datasets with ground truth are de-facto standards for evaluating post-hoc GNN explainers (for node classification), including GNNExplainer, PGExplainer, PGM-Explainer, SubgraphX, CF-GNNExplainer, RCExplainer, etc. (see Table 11 for more details); 2) without ground-truth information we can only compute some metrics, such as cosine similarity of explanation before and after attack (which will be defined later) to quantify attack performance. However, as shown in Table 1, the cosine similarity and the real attack performance (in terms of $\Delta\text{GEA}$, which will be defined later) are not necessarily related. Therefore, considering such metrics may not be meaningful.
4.2 Metrics and Baselines
@@
Graph Explanation Accuracy (GEA). We employ the graph explanation accuracy proposed in Agarwal et al. (2023) to quantify the quality of generated explanations when the ground-truth explanations are available. For simplicity, we assume that each edge is binary coded as ‘0’ (non-contributory) or ‘1’ (contributory) to model predictions in both ground-truth (provided by ShapeGGen) and predicted explanation masks (provided by the GNN explainer). Specifically, graph explanation accuracy measures the correctness of generated explanations by computing the Jaccard Index between ground-truth explanation $\mathbf{M}_{gt}$ and predicted explanation $\mathbf{M}_{pr}$ as
$\text{JAC}(\mathbf{M}_{gt},\mathbf{M}_{pr})=\frac{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})}{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FN}(\mathbf{M}_{gt},\mathbf{M}_{pr})},$ (14)
@@
where $\text{vec}(\cdot)$ flattens its input as a vector, and $\|\cdot\|_{F}$ represents the Frobenius Norm. Higher values of cos_sim (within $[-1,1]$) indicate higher graph explanation stability, and thus less successful attacks.
@@
Prediction Change. To measure the impact of edge perturbations on GNN predictions, we consider 1) $\Delta$Label: the ratio of nodes of which the predicted label has changed after attacks; and 2) $\Delta$Prob: the average absolute change of predicted probability (of the original predicted class) of all nodes.
@@
Baselines. Given the novelty of our problem setting, we only consider two trivial baselines for edge perturbations: 1) random flipping: we randomly flip $B_{1}$ edges; and 2) random rewiring: we randomly rewire $B_{2}$ edges within k-hop neighbours of the target node.
4.3 Attack Transferability Settings
@@
We consider attack transferability on the following four types of GNN explainers: 1) Gradient-based: Grad (Simonyan et al., 2013), GradCAM (Pope et al., 2019), GuidedBP (Baldassarre and Azizpour, 2019), Integrated Gradients (IG) (Sundararajan et al., 2017); 2) Perturbation-based: GNNExplainer (Ying et al., 2019), SubgraphX (Yuan et al., 2021); 3) Surrogate-based: PGMExplainer (Vu and Thai, 2020); and 4) Random Explainer that randomly generates random important edges.
For space reasons, all training protocols are deferred to Appendix G. Code is given in anonymous2024code.
5 Experimental Results and Analysis

### v2

@@ IV-B Relaxations and Loss Definitions
Relaxations. We assume that edge flipping graph $\mathcal{M}$ corresponding to $\mathbf{P}\in\{0,1\}^{n\times n}$ (namely using $\mathbf{P}$ as the adjacency matrix) is a Gilbert random graph. In other words, edge occurrences are conditionally independent from each other. As a result, with slight abuse of notations, the probability of graph $\mathcal{M}$ can be factorized as $P(\mathcal{M})=\prod_{1\leq i,j\leq n}P(m_{ij})$, dubbed $\tilde{\mathbf{P}}$. Next, we instantiate $P(m_{ij})$ using a Bernoulli distribution, namely $P(m_{ij})\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{ij})$ with $P(m_{ij}=1)=\tilde{\mathbf{P}}_{ij}$. As a result, Eq IV-A is equivalent to
@@
In this way, we relax the $\mathbf{P}\in\{0,1\}^{n\times n}$ from binary variables to continuous variables $\tilde{\mathbf{P}}\in[0,1]^{n\times n}$. Hence, we can utilise gradient-based methods to optimize the objective function. At the last step of the attack, we then discretize the result (see Section IV-C).
@@
$\underset{\tilde{\mathbf{P}}}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ (10)
$:=\underset{\tilde{\mathbf{P}}}{\max}~{}H(f_{\theta}(\mathbf{A},\mathbf{X}))-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ (11)
$\propto\underset{\tilde{\mathbf{P}}}{\max}~{}-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ (12)
$\approx\underset{\tilde{\mathbf{P}}}{\max}~{}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}})).$ (13)
Eq 11 holds by definition of mutual information; Eq 12 holds as $\theta,\mathbf{A},\mathbf{X}$ are fixed; Eq 13 holds when we replace conditional entropy $H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ with cross entropy $H(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ for practical optimization reason. Moreover, we define $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}+\tilde{\mathbf{P}}_{ij}$ if $\mathbf{A}_{ij}=0$, and $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}-\tilde{\mathbf{P}}_{ij}$ otherwise.
@@
$\underset{\tilde{\mathbf{P}}}{\min}~{}\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X})))$ (14)
$\propto\underset{\tilde{\mathbf{P}}}{\max}~{}\text{Dist}(\mathbf{M}_{s},\hat{\mathbf{M}}_{s})$ (15)
@@
where $\mathbf{M}_{s}$ and $\hat{\mathbf{M}}_{s}$ are the explanatory subgraphs on original graph and perturbed graph, respectively; $\mathbf{M}_{pr}$ and $\hat{\mathbf{M}}_{pr}$ are the predicted explanations (edge importance matrices) on original graph and perturbed graph, respectively. In other words, the explanatory subgraph is represented as an edge importance matrix in this study. Eq 15 holds by definition; Eq 16 holds when we define the $Dist$ function as the cosine distance metric, where $\text{vec}(\cdot)$ flattens a matrix to a vector, and $\|\cdot\|_{F}$ represents the Frobenius Norm.
Final Relaxed Objective. As a result, optimizing Eq IV-A amounts to optimizing
$\underset{\tilde{\mathbf{P}}\in[0,1]^{n\times n}\text{~{}s.t.~{}}\sum\tilde{\mathbf{P}}_{ij}<B}{\max}$ $\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}}))$
$+\beta*\left(1-\frac{\text{vec}(\mathbf{M}_{pr})\cdot\text{vec}(\hat{\mathbf{M}}_{pr})}{\|\mathbf{M}_{pr}\|_{F}\|\hat{\mathbf{M}}_{pr}\|_{F}}\right).$ (17)
As a result, $\mathcal{L}(\tilde{\mathbf{P}})$ (i.e., Eq IV-B) is differentiable w.r.t. $\tilde{\mathbf{P}}$.
IV-C Optimization
We employ Projected Gradient Ascent Algorithm (based on [46]) to optimize objective function (IV-B), and the details are given in Algorithm 1. To go back from the continuous $\tilde{\mathbf{P}}_{ij}$ to discrete ${\mathbf{P}}_{ij}$, we perform Bernoulli sampling with ${\mathbf{P}}_{ij}\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{ij})$ after optimization (Line 10). We note that the Projected Randomized Block Coordinate Descent (PR-BCD) [47] can perform this optimization more efficiently, but do not compare both methods due to the low scalability of the evaluated explainers.
@@ IV-C Optimization
10:    $\tilde{\mathbf{P}}_{t}\leftarrow\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})\leq B]}(\tilde{\mathbf{P}}_{t})$triangleright Project gradients, see Appendix Appendix A: Projected Gradient Descent for more detail
@@
IV-D Complexity Analysis
We first discuss the dense case here and then make a discussion of GNNs/explainers implemented with sparse matrices. The time complexity is as follows:
  1. 1.
  2. 2.
    Each Epoch:
    • Prediction and explanation on perturbed graph: $O(2T\cdot(f_{\theta}+g_{\lambda}))$, where $T$ is the number of epochs.
  3. 3.
    Sampling Step: $O(n^{2})$.
Meanwhile, the space complexity is as follows:
  1. 1.
  2. 2.
Assuming GNN and explainer are in $O(f_{\theta})=O(g_{\lambda})=O(Ed+nd^{2})$ with number of edges $E$, a dense attack would dominate the overall complexity ($O(n^{2})$). An immediate remedy would yield PRBCD [47] since its complexity is $O(B+E)$ with edge flipping budget $B$ and all additional components (e.g., loss function) could be implemented in $O(E)$ with sparse matrices. This yields an overall time and space complexity of $O(E)$, assuming $B\ll E$ and hidden dimensions $d\ll n$, which is the same asymptotic complexity as the sparse GNN $f_{\theta}$ and explainer $g_{\lambda}$.
V Experimental Setup
@@ V-A Datasets
To answer these questions, we need to evaluate the GNN explanation quality before and after the attacks. However, as pointed out by [50], evaluating the quality of post-hoc GNN explanations is challenging. This is because existing benchmark graph datasets lack ground-truth explanations or their explanations are not reliable. To mitigate this, they proposed a synthetic graph data generator ShapeGGen. It can generate a variety of graph datasets with ground-truth explanations, where graph size, node degree distributions, the ground-truth explanation sizes, and more can be varied. As a result, ShapeGGen allows us to mimic graph data in various real-world applications. More importantly, ShapeGGen can ensure that the generated data and its ground-truth explanations do not suffer from GNN explanation evaluation pitfalls [49], namely trivial explanations, redundant explanations, and weak GNN predictors. This makes ShapeGGen very suitable for studying the limitations of GNN explainers. We consider the synthetic datasets in Table I.
TABLE I: Summary of synthetic datasets with ground-truth explanations. #Nodes and #Edges represent the number of expected nodes and edges, respectively. We do not consider larger datasets for two reasons: 1) explanation accuracy will be very low ($<0.3$) for most explainers, and it is not meaningful to attack low accuracy explainers; 2) computation will be slow for the attacker (and the explainer).
Dataset Syn1 Syn2 Syn3 Syn4 Syn5 Syn6 Syn7
Shape House House House Circle House House House
#Subgraphs 10 50 50 50 50 50 100
Subgraph Size 6 6 10 10 10 10 10
P(Connection) 0.3 0.06 0.06 0.06 0.12 0.20 0.06
#Classes 2 2 2 2 2 2 2
#Nodes 59 299 521 511 489 492 1018
#Edges 164 970 1432 1314 1664 1898 3374
Average Degree 2.8 3.2 2.7 2.6 3.4 3.9 3.3
Max Budget 10 15 15 15 15 15 15
@@
Why not real-world datasets? We do not consider real-world datasets for two key reasons: 1) synthetic datasets with ground truth are de-facto standards for evaluating post-hoc GNN explainers (for node classification), including GNNExplainer, PGExplainer, PGM-Explainer, SubgraphX, CF-GNNExplainer, RCExplainer, etc. 2) without ground-truth information we can only compute some metrics, such as cosine similarity of explanation before and after attack (which will be defined later) to quantify attack performance. However, as shown in Table II, the cosine similarity and the real attack performance (in terms of $\Delta\text{GEA}$, which will be defined later) are not necessarily related. Therefore, considering such metrics may not be meaningful.
V-B Metrics and Baselines
@@
Graph Explanation Accuracy (GEA). We employ the graph explanation accuracy proposed in [50] to quantify the quality of generated explanations when the ground-truth explanations are available. For simplicity, we assume that each edge is binary coded as ‘0’ (non-contributory) or ‘1’ (contributory) to model predictions in both ground-truth (provided by ShapeGGen) and predicted explanation masks (provided by the GNN explainer). Specifically, graph explanation accuracy measures the correctness of generated explanations by computing the Jaccard Index between ground-truth explanation $\mathbf{M}_{gt}$ and predicted explanation $\mathbf{M}_{pr}$ as
$\text{JAC}(\mathbf{M}_{gt},\mathbf{M}_{pr})$
$=\frac{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})}{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FN}(\mathbf{M}_{gt},\mathbf{M}_{pr})},$ (18)
@@
where $\text{vec}(\cdot)$ flattens its input as a vector, and $\|\cdot\|_{F}$ represents the Frobenius Norm. Higher values of $\operatorname{Sim}_{\cos}$ (within $[-1,1]$) indicate higher graph explanation stability, and thus less successful attacks.
@@
Prediction Change. To measure the impact of edge perturbations on GNN predictions, we consider two metrics as follows: 1) $\Delta$Label: the ratio of nodes of which the predicted label has changed after attacks; and 2) $\Delta$Prob: the average absolute change of predicted probability (of the original predicted class) of all nodes, which is formally defines as
$\Delta\text{Prob}=\frac{\sum_{i=1}^{N}\lvert P(y_{i}=y_{i}^{*})-\hat{P}(y_{i}=y_{i}^{*})\rvert}{N},$
where $P(y_{i}=y_{i}^{*})$ is the prediction probability of node $v_{i}$ being classified as $y_{i}^{*}$ (the original predicted class with the highest probability) on the original graph. Moreover, $\hat{P}(y_{i}=y_{i}^{*})$ is the prediction probability of node $v_{i}$ being classified as $y_{i}^{*}$ (the original predicted class with the highest probability on the original graph) on the perturbed graph.
@@
Attacker Baselines. Given the novelty of our problem setting, we only consider two trivial baselines for edge perturbations: 1) random flipping: we randomly flip $B_{1}$ edges; and 2) random rewiring: we randomly rewire $B_{2}$ edges within k-hop neighbours of the target node.
V-C Attack Transferability Settings
@@
We consider attack transferability on the following five types of GNN explainers: 1) Gradient-based: GradCAM [13], Integrated Gradients (IG) [32]; 2) Perturbation-based: GNNExplainer [11], SubgraphX [17]; 3) Surrogate-based: PGMExplainer [12]; 4) Decomposition-based methods: GNN-LRP [34] and 5) Random Explainer that randomly generates random important edges.
V-D Training Protocols and Compute Resources
GNN Predictor Training Protocols. To avoid pitfalls caused by weak GNN predictor [49], we should train the GNN predictor close to its maximum possible performance. When training the GNN predictor, we employ an Adam optimizer where the learning rate is 1e-2, weight decay is 1e-5, the number of epochs is 300, and the hidden dimension is 32.
GNN Explainer Training Protocols. We followed the authors’ recommendations for setting the hyperparameters of GNN explanation methods. We select top-$k~{}(k=25\%)$ important edges to generate explanations for all GNN explainers. Particularly, when training the PGExplainer, the number of training epochs is 10.
GNN Explanation Attacker Training Protocols. We set the trade-off hyperparameter in the final loss as $0.1$ such that the magnitudes of the two loss terms are approximately the same. Besides, the learning rate in gradient ascent is set to $0.1$. Importantly, we set the maximal training epochs as 100 due to computation time reason, while further increasing this number can largely improve the attack performance (in terms of $\Delta$GEA). Moreover, we set the perturbation budget to 15 at maximum for synthetic data (note that we attempt to employ a hard minimum budget of 5 with 2000 maximum trials of Bernoulli sampling). We perform sensitivity analysis regarding these hyperparameters, and the results and analysis will be discussed in Section VI-D.
Explanation Accuracy Computation. The $get\_acc(\cdot)$ function of GraphXAI considers all non-zero edges as predicted positives (namely with binary code ‘1’) by default when computing JAC. Following [50], we set the top 25% (of edges with the highest edge importance scores) as 1 and the rest as zeros in our evaluations when computing both GEA and cosine similarity.
Software and Hardware. The evaluated GNN explainers are implemented mainly based on GraphXAI  [^1] https://github.com/mims-harvard/GraphXAI [50]. Algorithms are implemented in Python 3.8 (using PyTorch [51] and PyTorch Geometric [52] libraries when applicable) and ran on internal clusters with AMD 128 EPYC7702 CPUs (with 512GB RAM). All experiments can be finished within 7 days if using 10 such machine instances. All code and datasets are available on GitHub [^2] https://github.com/ZhongLIFR/GXAttack.
VI Experimental Results and Analysis